From GlueXWiki
Jump to: navigation, search

November 10, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production update
    • Second package tests (Lubomir)
  2. Engineering update; card positions (Bill)
  3. Electronics update (Chris); Fernado's LV scheme
  4. Test set-up at EEL126 OSP,THA, [(Beni)
  5. TDR (Lubomir)
  6. Other: new FDC logbook


Participants: Eugene, Bill, Dave, Chris, Simon, Vladimir, Beni, and Lubomir.


- Dave on second package production: first cell was installed, all the elements for the second cell are ready except to populate the wire boards (Chris will do it tomorrow). Cathodes for the third cell also ready, working on the fourth cell cathodes. Third wire plane was strung and wires glued and soldered. The HV caps were replaced on the fourth, but we have problems so far with the dark current there; may require additional change of some capacitors. The fifth wire frame was laminated today. Problems with field wire spools, Dave will discuss it with Steve Christo from Hall B.

- Lubomir: The installation of the first cell was much easier due to the new hubs/pins Bill designed. Flushing with gas since Tuesday, chamber is very tight (bubbles at as low as 15 ccpm). Setting HV on it was also easy. Vladimir tested all the channels with a scope and 55Fe: all working fine. But again, we see a noise at the same place (around HV trace on the fourth +HV sector) we saw on three cells in the first package, this time with higher frequency (~2.5 kHz). Will rotate the top cathode by 30deg after the meeting to make sure we have the same problem here.


- The scheme for the positions of the signal connectors in the package is linked above

- Bill: will look into the cooling system. Using Fluorinert (FC-72 made by 3M) instead of water require proper sealing. We discussed if we will need real cooling for the 126 tests, or just the tubing.


- Chris is ready with the first 22 pre-amps that handed to us at the meeting. Cody will test the rest with a LabView system; working on the program now. Will install the pre-amps and the first cooling tube on the first package tomorrow.

- Based on the Bill's drawing of the pre-amp cards, Fernanda made a scheme for the low voltage cables, also linked above. Will have 4 cards per cable.

Chamber testing at 126

- Beni: working on the OSB and THA for the tests at 126 (latest version linked above). Some leakage on an Indiana chamber, made temporary fix. The materials for the translator boards are here, now writing PR for their assembly.


- Slow progress, current version is at the link above. Need drawings from Bill of the wire and cathode frames. Simon working on the tracking section. Eugene: the document should be self contained as much as possible, i.e. important info must be in the document not cited.


- Beni setup a new logbook to be used for the FDC (see link above).